Mr Tan et off from Town A to Town B at an average speed of 75km/h at 08 00. Mr Chen set off from Town A after some time and overtook Mr Tan at 10 00. When Mr Chen reached Town B at 12 00, Mr Tan was still 30km away.
(a)What was Mr Chen's average speed?
(b)What time did Mr Chen leave Town A?
pls. try not to use algebra but if there is no choice, then, algebra would be fine.
: Mr Tan et off from Town A to Town B at an average speed of 75km/h at 08 00. Mr Chen set off from Town A after some time and overtook Mr Tan at 10 00. When Mr Chen reached Town B at 12 00, Mr Tan was still 30km away.
(a)What was Mr Chen's average speed?
(b)What time did Mr Chen leave Town A?
pls. try not to use algebra but if there is no choice, then, algebra would be fine.

Answer by vleith(1174)

Not sure how one would solve this without using algebra:) But will try to get it simple.
Tan left at 8 going 75. Chen caught up to him at 10. So when Chen passes Tan, Tan has been driving for 2 hours. He has gone 75*2 = 150km.
Chen reaches B in another 2 hours. But Tan is still 30km from town. Tan has driven another 2 hours since being passed by Chen, so Tan has driven another 150km. And he is till 30km from B. So the total distance between A and B is 150+150+30 = 330km.
1) The distance from when Chen passed Tan to B is 150+30 = 180.
Chen took 2 hours to cover 180km, so Chen's average speed is 180/2 = 90km.
2) The distance from A to where Chen catches Tan is 150km. Chen's average speed is 90. 150/90 = 5/3 hours. Chen left 20 minutes after Tan. So Chen left at 8:20